Find the mean, variance, and standard deviation for the number of heads when 15 coins are tossed.
15(2)=30

Answer by oscargut(2103)

mean = np = 15(0.5) = 7.5
variance = np(1-p) = 15(0.5)(0.5) = 3.75
If you need more help today, you can contact me now at:
mthman@gmail.com
Thanks